OpenStudy (samigupta8):

The number of real x satisfying the equation (x-2)[x]={x}-1

OpenStudy (samigupta8):

I m getting only 1 value of x which is 1

OpenStudy (samigupta8):

@priyar @hartnn @ganeshie8 @wio @michele_laino

OpenStudy (samigupta8):

Here ,{x} stands for fractional part function And [x] is greatest integer function

Parth (parthkohli):

\(x=1\) is obviously a solution. Other than that, we know that the right hand side is negative and lies in \([-1,0)\). So \(x-2\) and \([x]\) have opposite signs. That restricts our domain to \([0,2]\). \([x]\) is obviously not zero. \([x]\) can be one. We've already found one such solution. Let's try to find others. \(x-2 = \{x \} -1 \Rightarrow \lfloor x\rfloor =1\). This already tells us that there are infinite solutions.

Parth (parthkohli):

All reals in the interval \([1, 2)\) satisfy that, so the answer is infinity.

OpenStudy (samigupta8):

Okk Shall i tell you my way now..which strictly bans over [x] to be equal to 1

OpenStudy (samigupta8):

We can write x as [x]+{x} Now lhs would become [x]^2+[x]{x}-2[x] Equating , we will be left with (1-[x])^2={x}(1-[x]) Now , {x}= (1-[x])^2/1-[x]

OpenStudy (samigupta8):

Now this clearly tells that [x] can never be equal to 1

OpenStudy (priyar):

Well, i am getting x belongs to [1,infinity)..is it correct?

Parth (parthkohli):

It's not.

OpenStudy (samigupta8):

Yep !!..he's correct...

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

do u need help

OpenStudy (samigupta8):

Yep ! Will you help?

OpenStudy (anonymous):

sure

OpenStudy (samigupta8):

Do so then!

Parth (parthkohli):

(x - 2)[x] = {x} - 1 x = 10 8*10 = 0 - 1? The answer is not 1 to infinity...

OpenStudy (samigupta8):

@parthkohli what is the fault in method?

OpenStudy (samigupta8):

My method*

Parth (parthkohli):

The fault is that you get the wrong solution. =P

OpenStudy (samigupta8):

Like seriously!! What is that thing in my method which is resulting in that fault?

Parth (parthkohli):

In the second-last line, you wrote this:\[(1-|x|)^2 = \{ x \}(1 - |x|)\]This equation says that whenever \(|x| = 1\), this equation will be satisfied. This itself tells us that the equation will have infinite solutions.

OpenStudy (samigupta8):

Bt if u solve this to take the value for {x} U get {x}=(1-[x])^2/1-[x]

OpenStudy (samigupta8):

and for it to hold true we must say that [x] is not equal to 1

Parth (parthkohli):

That's very illegal in maths. You just changed the equation to something that it's not.\[(1-x)^2 = k(1-x)\]Are you saying that \(x = 1\) does not satisfy the above equation just because if we solve for \(k\)\[k = \frac{(1-x)^2}{1-x}\]it does not allow \(x = 1\)? But \(x = 1\) does satisfy the original equation and the original equation is all we care about.

OpenStudy (samigupta8):

Okay!!

OpenStudy (samigupta8):

Ty @parthkohli

Parth (parthkohli):

<3

OpenStudy (michele_laino):

if I add \([x]\) to both side, I get: \[\Large \begin{gathered} \left( {x - 2} \right)\left[ x \right] + \left[ x \right] = \left\{ x \right\} - 1 + \left[ x \right] \hfill \\ \hfill \\ \left( {x - 1} \right)\left[ x \right] = x - 1 \hfill \\ \end{gathered} \] being: \[\Large \left\{ x \right\} + \left[ x \right] = x\]